how would i rewrite a quotient as multiplication by the reciprocal here is an example. 12/(1/2)
Answer by stanbon(75887)

how would i rewrite a quotient as multiplication by the reciprocal here is an example. 12/(1/2)
-----
[12/(1/2)][(2/1)/(2/1)] = (12*2)/(1/2*2/1) = 24/1 = 24
========================================================
cheers,
Stan H.